RB Music Store assembles stereo equipment for resale in the shop. The store offers two products; two products; turntables and cassette players. The store makes a profit of P400 on each turntable and P250 on each cassette. Both must go through two steps-assembly and bench checking. A turntable takes qw hours to assemble and 4 hours to bench check. A cassette player takes 4 hours to assemble but 8 hours to bench check. Looking at this month’s schedule, the store sees that it has 60 assembly hours uncommitted and 40 hours bench-cutting time available. What combination of turntables and cassette players would maximize the store’s profit for the month?


1
Expert's answer
2022-04-11T15:08:07-0400

Let "x_1" - number of turntables, and "x_2" - number of cassette players.

"\phi=400x_1+250x_2 \to max"

"\begin{cases}
14x_1+4x_2\leq60\\
4x_1+4x_2\leq40
\end{cases}"

Let 's build graphs of the resulting equations:



The answer to our problem will be at point C:

"\begin{cases}
14x_1+4x_2=60\\
4x_1+4x_2=40
\end{cases}\implies\begin{cases}
x_1=2\\x_2 = 8
\end{cases}"

"\phi=400*2+8*250=2800"

Answer: 2 turntables and 8 cassette players.